OpenStudy (mrchair):

Find the sum. (13x2 + 5x) + 8x2 = 26x2 26x5 20x2 + 5x 21x2 + 5x

OpenStudy (anonymous):

21x^2 + 5x

OpenStudy (anonymous):

lol @Ash_Uchiha

OpenStudy (anonymous):

:) beat cha that time! @shayforrlyfe

OpenStudy (anonymous):

\[(13x^2+5x)+8x^2\] \[169+5x+8x^2\] \[-169 = -169\] \[\frac{ 5 }{ 5 } = \frac{ -105 }{ 5 }\] \[= -21\]

OpenStudy (anonymous):

aha yea you did xD

OpenStudy (mrchair):

i already finished this on my own lol i got a new one!

OpenStudy (mrchair):

thats wrongs freaks not an option

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

LOL

OpenStudy (mrchair):

don't be meen XD

OpenStudy (anonymous):

We're not XD